OpenStudy (anonymous):

2-3/4x=20 solve for the equation

OpenStudy (anonymous):

-3/4x=20 x=-15

OpenStudy (anonymous):

x=-24

OpenStudy (anonymous):

no, it's like 2-0.75x=20

OpenStudy (anonymous):

the question was 2-3/4x=20

OpenStudy (anonymous):

the other way is: 1/4x=20 x=80

OpenStudy (anonymous):

or is it 1/4x=20 x=1/80 ?

OpenStudy (anonymous):

Wrong. -3/4x=20-2 -3/4x=18 -3x=18*4 -3x=72 x=72/-3 x=-24

OpenStudy (anonymous):

The monster ate the cookie :P

OpenStudy (anonymous):

wrong to you step two to 3: -3/4x=18 / -3x=72 ^

OpenStudy (anonymous):

How is that wrong?

OpenStudy (anonymous):

@x3pufflix3:I don't see anything wrong in your answer! :)

OpenStudy (anonymous):

-3/x=72 1/x=-24 x=-1/24

OpenStudy (anonymous):

he missed dividing

OpenStudy (anonymous):

or she

OpenStudy (anonymous):

he is correct use the check method to veify

OpenStudy (anonymous):

Blueyedgem I gave you my answer its your choice to choose whose right and wrong. I donot like to argue with anyone. So ciao buddy! ;D

OpenStudy (anonymous):

Wait she assumes \[\frac{-3}{4}x\] so did I